OpenStudy (anonymous):

derivative: (x-1)^(x)^1/2

OpenStudy (anonymous):

\[(x-1)^{\sqrt{x}}\]

OpenStudy (anonymous):

i took the log and got ln(x-1)\[\sqrt{x}\]

OpenStudy (anonymous):

\[\ln(x-1)\sqrt{x}\]

OpenStudy (experimentx):

differentiate using product rule.

OpenStudy (anonymous):

Then i used the product rule and got\[(\sqrt{x}/(x-1)) + (\ln(x-1)/2\sqrt{x)}\]

OpenStudy (anonymous):

is my answer correct?

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

Also, how do you know when to use the quotient rule? sometimes given a fraction, we are asked to solve it with the product rule.

OpenStudy (anonymous):

thank you :)

OpenStudy (experimentx):

quotient rule is same as product rule u/v = uv^-1

OpenStudy (experimentx):

and you are missing something.

OpenStudy (anonymous):

oh ok and for limits when do we use the L'hospital's rule?

OpenStudy (anonymous):

what am i missing?

OpenStudy (anonymous):

L'hospital's is used when the limit you are evaluating is indeterminate, such as 4/0 or 0/0. You don't take the quotient rule with L'hospital' though, you take the derivative of the numerator and denominator separately

OpenStudy (anonymous):

@brainshot3 thank you!

OpenStudy (experimentx):

lny = x something, <---- took derivative here 1/y dy/dx = x <--- derivative of x's dy/dx = y times derivative of x's

OpenStudy (anonymous):

is this implicit differentiation? i don't really understand it

OpenStudy (experimentx):

it's same as differentiation, but you will have two variables instead of one.

OpenStudy (anonymous):

oh ok

OpenStudy (experimentx):

check mit ocw video on single variable calculus.

OpenStudy (anonymous):

so my answer should be \[lny=\sqrt{x}/x-1 + \ln(x-1)/2\sqrt{x}\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (experimentx):

no your answer should be y times the value of left. and put the value of y.

OpenStudy (anonymous):

You want dy/dx alone on one side and you do that by multiplying out the y

OpenStudy (anonymous):

\[dy/dx=y \sqrt{x}/x-1 + \ln(x-1)/2\sqrt{x} \]

OpenStudy (anonymous):

Looks correct

OpenStudy (anonymous):

because the derivative of ln y = 1/y

OpenStudy (anonymous):

thank you all for your help :)

OpenStudy (anonymous):

Wait, you forgot to distribute the y

OpenStudy (experimentx):

y = (x-1)^x so your answer should be \[dy/dx = (x-1)^x (√x/(x−1)+ln(x−1)/2√x) \]

OpenStudy (anonymous):

oh i got it now! thank you :)